# Physics-informed Kolmogorov–Arnold Network with Chebyshev Polynomials for Fluid Mechanics

### v2

@@ III.6 Two-dimensional Steady Navier–Stokes Equation
This section studies the lid-driven cavity flow at $Re$ = 3200 to assess ChebPIKAN’s capability in approximating steady flows. Since the temporal derivative term can be disregarded, the governing equations simplify to
$u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}=-\frac{\partial p}{\partial x}+\nu\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right),$ (37)
$u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}=-\frac{\partial p}{\partial y}+\nu\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right).$ (38)
@@
$\begin{split}Los{s_{PDE}}=\,&u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}+\frac{\partial p}{\partial x}-\nu\left(\frac{\partial^{2}u}{\partial x^{2}}+\frac{\partial^{2}u}{\partial y^{2}}\right)\\ +\,&u\frac{\partial v}{\partial x}+v\frac{\partial v}{\partial y}+\frac{\partial p}{\partial y}-\nu\left(\frac{\partial^{2}v}{\partial x^{2}}+\frac{\partial^{2}v}{\partial y^{2}}\right)\\ +&\frac{\partial u}{\partial x}+\frac{\partial v}{\partial y}.\end{split}$ (40)
@@
Training ANN 1.80×$\times 10^{-2}$ 1.40×$\times 10^{-3}$ 2.81×$\times 10^{-4}$ 1.17×$\times 10^{-4}$
KAN 1.15×$\times 10^{-3}$ 5.66×$\times 10^{-5}$ 9.24×$\times 10^{-6}$ 5.00×$\times 10^{-6}$
PINN 3.56×$\times 10^{-2}$ 7.33×$\times 10^{-3}$ 1.45×$\times 10^{-3}$ 1.17×$\times 10^{-3}$
ChebPIKAN 6.02×$\times 10^{-3}$ 4.19×$\times 10^{-4}$ 7.61×$\times 10^{-5}$ 6.89×$\times 10^{-5}$
PDE PINN 3.51×$\times 10^{-2}$ 7.36×$\times 10^{-3}$ 4.33×$\times 10^{-3}$ 4.52×$\times 10^{-3}$
ChebPIKAN 6.22×$\times 10^{-3}$ 7.55×$\times 10^{-4}$ 1.38×$\times 10^{-4}$ 8.32×$\times 10^{-5}$
Testing ANN 3.85×$\times 10^{-2}$ 2.41×$\times 10^{-2}$ 2.96×$\times 10^{-2}$ 9.41×$\times 10^{-3}$
KAN 1.48×$\times 10^{-2}$ 2.49×$\times 10^{-2}$ 2.20×$\times 10^{-2}$ 2.92×$\times 10^{-2}$
PINN 5.88×$\times 10^{-3}$ 8.02×$\times 10^{-3}$ 1.84×$\times 10^{-3}$ 1.45×$\times 10^{-3}$
ChebPIKAN 1.15×$\times 10^{-2}$ 2.31×$\times 10^{-3}$ 8.42×$\times 10^{-3}$ 4.45×$\times 10^{-3}$
@@
As evidenced in Figure 22 (b), PINNs and ChebPIKANs exhibit superior performance on unseen test data compared to standard models. Notably, while KANs achieve the best results on the known training dataset, their performance on the test set is inferior even to PINNs. This suggests that for physical problems where only sparse real-world data is available for training, physics-informed frameworks can leverage minimal ground-truth data as anchor points and reconstruct the entire flow field using governing equations. Such an approach effectively mitigates overfitting and enhances the overall predictive capability over the entire field. Furthermore, the four-hidden-layer KANs, despite delivering optimal training accuracy, perform the worst on the test set. This not only highlights the critical challenge of overfitting in practical applications but also underscores that neural networks with strong fitting capacities—when faced with insufficient data—require carefully designed physics-informed loss functions to improve their real-world performance.
@@ III.7 The Impact of the Number of Trainable Parameters on Neural Network Performance
30 9.13×$\times 10^{-2}$ 6.21×$\times 10^{-2}$ 4.04×$\times 10^{-2}$
55 4.02×$\times 10^{-2}$ 3.56×$\times 10^{-2}$ 1.97×$\times 10^{-2}$
80 3.36×$\times 10^{-2}$ 3.51×$\times 10^{-2}$ 2.56×$\times 10^{-2}$
105 3.52×$\times 10^{-2}$ 2.85×$\times 10^{-2}$ 1.93×$\times 10^{-2}$
@@
38 2.06×$\times 10^{-1}$ 2.43×$\times 10^{-1}$ 7.27×$\times 10^{-2}$
68 1.58×$\times 10^{-1}$ 1.22×$\times 10^{-1}$ 7.77×$\times 10^{-2}$
98 2.85×$\times 10^{-1}$ 2.73×$\times 10^{-1}$ 1.02×$\times 10^{-1}$
128 1.60×$\times 10^{-1}$ 2.77×$\times 10^{-1}$ 7.46×$\times 10^{-2}$
@@
III.8 The Performance of Flow Field Prediction in Temporal Extrapolation
@@ IV Conclusions and Discussion
Chebyshev physics-informed Kolmogorov–Arnold network (ChebPIKAN) which incorporates physical constraints into KAN through PDE loss terms while replacing the Kolmogorov-Arnold theorem with Chebyshev polynomial, achieves a transformation comparable to the evolution from conventional ANNs to PINNs. In this study, the performance improvement of ChebPIKAN with added physical information relative to KAN was verified by solving partial differential equations, and ChebPIKAN has obvious advantages. It is reflected in the higher stability and generalization of the network compared to KAN without adding physical information.
@@
Since the localization of KAN fitting ability, ChebPIKAN has better generalization performance. This advantage is evident in the continuous learning performance of the network and its resilience during debugging and training across various tasks, which effectively mitigate the risk of catastrophic forgetting. Furthermore, when future tasks involve overlapping or similar samples, the KAN structure is poised to offer significant benefits.

### v3

@@ III.6 Two-dimensional Steady Navier–Stokes Equation
This section studies the lid-driven cavity flow at $Re$ = 3200 to assess ChebPIKAN’s capability in approximating steady flows. Since the temporal derivative term can be disregarded, the governing equations are simplified to
@@
[…]
@@
As evidenced in Figure 22 (c), PINNs and ChebPIKANs exhibit superior performance on unseen test data compared to standard models. Notably, while KANs achieve the best results on the known training dataset, their performance on the test set is inferior even to PINNs. This suggests that for physical problems where only sparse real-world data is available for training, physics-informed frameworks can leverage minimal ground-truth data as anchor points and reconstruct the entire flow field using governing equations. Such an approach effectively mitigates overfitting and enhances the overall predictive capability over the entire field. Furthermore, the four-hidden-layer KANs, despite delivering optimal training accuracy, perform the worst on the test set. This not only highlights the critical challenge of overfitting in practical applications but also underscores that neural networks with strong fitting capacities—when faced with insufficient data—require carefully designed physics-informed loss functions to improve their real-world performance.
@@ III.7 The Impact of the Number of Trainable Parameters on Neural Network Performance
[…]
@@
[…]
@@
III.8 Flow Field Prediction Performance under Varied Reynolds Number Conditions
The generalization capability across different Reynolds numbers is crucial for validating the performance of machine learning methods in fluid prediction problems. We systematically evaluate the ChebPIKAN architecture comprising four hidden layers of 30 neurons each, analyzing its flow field prediction capability across four Reynolds number regimes as quantified in Table 11.
Table 11: Mean residuals of ChebPIKAN for steady Navier–Stokes equation across different Reynolds numbers.
Re Hidden Layers Er(u) Er(v) Er(p)
400 4x30 0.09% 0.19% 0.25%
1000 4x30 0.20% 0.24% 0.17%
3200 4x30 0.13% 0.14% 0.10%
5000 4x30 0.29% 0.21% 0.16%
Table 11 demonstrates that when ChebPIKAN contains sufficient trainable parameters to represent physical flow characteristics, it maintains prediction errors below 0.3% across sub-5000 Reynolds number cases despite minor performance fluctuations. This confirms the method’s generalizability within this Reynolds number range.
The current implementation of ChebPIKAN shares a common limitation with most contemporary intelligent fluid mechanics methodologies, requiring retraining when applied to different Reynolds number regimes. To achieve the ultimate goal of replacing conventional CFD approaches, future research should focus on developing either rapid transfer learning techniques or alternative architectures that directly incorporate the Reynolds number as an input variable. These advancements will be critical for establishing truly generalizable flow prediction systems.
III.9 The Performance of Flow Field Prediction in Temporal Extrapolation
@@ IV Conclusions and Discussion
Chebyshev physics-informed Kolmogorov–Arnold network (ChebPIKAN) which incorporates physical constraints into KAN through PDE loss terms while replacing the Kolmogorov-Arnold theorem with Chebyshev polynomials, achieves a transformation comparable to the evolution from conventional ANNs to PINNs. In this study, the performance improvement of ChebPIKAN with added physical information relative to KAN was verified by solving partial differential equations, and ChebPIKAN has obvious advantages. It is reflected in the higher stability and generalization of the network compared to KAN without adding physical information.
@@
Since the local fitting capability of KAN, ChebPIKAN has better generalization performance. This advantage is evident in the continuous learning performance of the network and its resilience during debugging and training across various tasks, which effectively mitigate the risk of catastrophic forgetting. Furthermore, when future tasks involve overlapping or similar samples, the KAN structure is poised to offer significant benefits.
@@ Acknowledgements.
Appendix A Performance Observations of the Original KAN
Following its introduction, the Kolmogorov-Arnold Network (KAN) has garnered significant attention in the scientific community, prompting our research team to promptly engage in its investigation. However, during our application of the original KAN framework to flow field prediction tasks, we consistently observed prediction inaccuracies occurring at domain boundaries, as illustrated in Figure A1.
Figure A1: Phenomena observed during flow field prediction using the original KAN.
Figure A1 clearly shows significant errors occurring at the boundaries of the computational domain when using the original KAN for flow prediction, which led us to investigate the causes of this phenomenon. From the mathematical perspective, B-splinessplines exhibit potential limitations when applied to the complex problems investigated in this study. Primarily, their non-orthogonal basis functions exhibit linear dependence, which may lead to solution instability when strong inter-basis correlations occur. Second, while spline curves are commonly used in research to prevent Runge’s phenomenon, they still exhibit severe oscillations when processing highly irregular node distributions. In contrast, Chebyshev polynomials inherently overcome these limitations through their dual advantages of mutual orthogonality and endpoint-clustering characteristics, prompting us to incorporate them into the original KAN framework.
Appendix B Prediction Performance under Different Weighting Coefficients $\lambda$
For neural networks with physical information loss terms, the weighting coefficient $\lambda$ determines the relative contribution of $Los{s_{data}}$ and $Los{s_{phy}}$ in guiding the optimization direction during training, and remains a critical hyperparameter affecting model performance. In the main study, we first conduct pretraining with $\lambda$ = 1, then determine the final weighting coefficient $\lambda$ = 0.1 based on the ratio between $Los{s_{data}}$ and $Los{s_{phy}}$.
In this Appendix B, we investigate the impact of different $\lambda$ values on neural network performance. We evaluate both ChebPIKAN and PINN architectures with 2×30 hidden layers using $\lambda$ values ranging from $10^{-4}$ to $10^{2}$, with their predictive performance summarized in the Table B1.
Table B1: Mean residuals of ChebPIKANs and PINNs for Navier–Stokes equations under different weighting coefficients $\lambda$.
Name Hidden Layers $\lambda$ Er(u) Er(v) Er(p)
ChebPIKANs 2x30 $10^{-4}$ 3.05% 0.75% 3.70%
2x30 $10^{-3}$ 1.78% 1.05% 2.01%
2x30 $10^{-2}$ 0.44% 0.38% 0.49%
2x30 $10^{-1}$ 0.24% 0.35% 0.33%
2x30 $10^{0}$ 1.09% 1.39% 1.16%
2x30 $10^{1}$ 8.86% 12.40% 16.80%
2x30 $10^{2}$ 29.15% 44.96% 54.03%
PINNs 2x30 $10^{-4}$ 2.41% 2.73% 4.31%
2x30 $10^{-3}$ 1.93% 2.82% 3.89%
2x30 $10^{-2}$ 2.32% 2.72% 3.93%
2x30 $10^{-1}$ 2.10% 2.52% 2.34%
2x30 $10^{0}$ 7.65% 8.60% 5.53%
2x30 $10^{-1}$ 21.83% 32.97% 40.04%
2x30 $10^{-2}$ 37.07% 54.99% 57.91%
Figure B1: Mean residuals of $u$, $v$, and $p$ of ChebPIKANs and PINNs with different weighting coefficients $\lambda$ for Navier–Stokes equations. (a) ChebPIKANs, (b) PINNs.
As shown in Table B1 and Figure B1, ChebPIKANs consistently demonstrate superior performance to PINNs under various $\lambda$ conditions. Both architectures achieve their best performance at $\lambda$ = 0.1 among the discrete values we examined, suggesting that the optimal $\lambda$ probably lies around this value. Therefore, we select $\lambda$ = 0.1 in our research. The results demonstrate that network performance degrades significantly when the weighting coefficient $\lambda$ becomes either too small or too large, revealing an inherent trade-off in balancing the two loss terms. At lower $\lambda$ values, the $Los{s_{data}}$ dominates the optimization process, causing the network to behave similarly to conventional KANs and ANNs without physical constraints. Conversely, higher $\lambda$ values may drive the training process away from the training data distribution as the model prioritizes minimizing the $Los{s_{phy}}$, potentially leading to constant-field solutions that satisfy the physical constraints while degrading the prediction accuracy.
This analysis highlights how the pretraining process with limited $\lambda$ variations may fail to identify optimal weighting coefficients, since the initial choice of $\lambda$ itself influences the final parameter selection. The challenge of determining appropriate hyperparameter values remains a persistent issue in physics-informed neural network development. Consequently, the creation of adaptive algorithms that can automatically balance these competing loss terms during training represents a valuable direction for future research in this field. We will continue to develop our algorithm specifically to address this issue.